State the equations of the verticaland horizontal asymptotes of g(x)= 8/x+5 +6
Answer by stanbon(75887)

State the equations of the vertical and horizontal asymptotes of
g(x)= 8/(x+5) +6
----
Vertical when x+5 = 0 ;; x = -5
----------------------------------
Horizontal when y = 0*x/1*x = 0
=========================
Cheers,
Stan H.
==========================